# LSMS: Language-guided Scale-aware MedSegmentor for Medical Image Referring Segmentation

### v2

@@ 4.6 Visualization Analysis
In Figure 5, we illustrate segmentation results and feature maps obtained from two sets of inputs, denoted as (a) and (b). In Figure 5(a), the language expression is “An irregular mass in the left lobe of the liver, with pronounced inhomogeneity.” The language expression for Figure 5(b) is “The smaller low-density nodule, 27 mm in diameter.” The labels $S_{i},i\in\{1,2,3,4\}$ represent encoded multi-modal features from different stages. The segmentation results demonstrate LSMS’s ability to accurately locate and segment the specified lesions based on the language expressions. Analyzing $S_{i},i\in\{1,2,3,4\}$ reveals LSMS’s progressive focus from shallow to deep layers onto the corresponding lesions: $S_{1}$ comprehends the overall visual context, $S_{2}$ extensively attends to various objects within the image, $S_{3}$ narrows down to candidate lesions, and $S_{4}$ precisely focuses on the lesion specified by the language input. In sample (a), where only one large lesion is present, $S_{3}$ rapidly focuses on the target area. As modeling deepens, $S_{4}$ demonstrates profound visual-linguistic cues. In sample (b), where multiple lesions coexist in the image, $S_{3}$ exhibits multiple attention points. Through further vision-language interaction, $S_{4}$ can focus on the lesions specified by the expression, aiding the model in making correct predictions.
5 Conclusion
@@
In this paper, we introduce a novel task, Medical Image Referring Segmentation, driven by clinical demands. And we propose the LSMS model with two appealing designs. LSMS integrates a novel attention mechanism to enhance object localization by tightly interacting scale-aware visual knowledge and linguistic cues. We introduce a full-scale decoder for global modeling of multi-modal features, improving boundary prediction in segmentation. Additionally, we develop the RefHL-Seg dataset for training and validating the MIRS task. Our experimental results show LSMS outperforms current methods in both MIRS and conventional medical image segmentation tasks with lower computational costs.

### v3

@@ IV-F Visualization Analysis
In Fig. 6, we illustrate segmentation results and feature maps obtained from two pairs of inputs, denoted as (a) and (b). In Fig. 6(a), the language expression is “An irregular mass in the left lobe of the liver, with pronounced inhomogeneity.” The language expression for Fig. 6(b) is “The smaller low-density nodule, 27 mm in diameter.” The labels $P_{i},i\in\{1,2,3,4\}$ represent encoded multi-modal features from different stages. The segmentation results demonstrate LSMS’s ability to accurately locate and segment the specified lesions based on the language expressions. Analyzing $P_{i},i\in\{1,2,3,4\}$ reveals LSMS’s progressive focus from shallow to deep layers onto the corresponding lesions: $P_{1}$ comprehends the overall visual context, $P_{2}$ extensively attends to various objects within the image, $P_{3}$ narrows down to candidate lesions, and $P_{4}$ precisely focuses on the lesion specified by the language input. In sample (a), where only one large lesion is present, $P_{3}$ rapidly focuses on the target area. As modeling deepens, $P_{4}$ demonstrates profound visual-linguistic cues. In sample (b), where multiple lesions coexist in the image, $P_{3}$ exhibits multiple attention points. Through further vision-language interaction, $P_{4}$ can focus on the lesions specified by the expression, aiding the model in making correct predictions.
V Conclusion
@@
In this paper, we introduce a new task, Referring Lesion Segmentation, driven by clinical demands. To support this task, we develop the RefHL-Seg benchmark for training and validating RLS. We propose LSMS for lesion segmentation with two appealing designs. LSMS integrates a novel attention mechanism to enhance object localization by tightly interacting scale-aware visual knowledge and linguistic cues. We introduce a full-scale decoder for global modeling of multi-modal features, improving boundary prediction in segmentation. Additionally, we design a specialized loss function to enhance fine-grained discrimination. Our experimental results show LSMS outperforms existing methods in both RLS and conventional lesion segmentation tasks with lower computational costs.